# bibliography.bib

### v1

@@ 1 Introduction
Figure 1: Example of safe paths and safe sequences for the $s$-$t$ path covers in a DAG $G$ with unique source $s$ and unique sink $t$. That is, for every path cover of $G$ (i.e. set of paths from $s$ to $t$ in $G$ such that every arc of $G$ appears in a path of the cover), there is a path in the cover containing the safe path or the safe sequence. We can see that safe sequences extend safe paths over complex subgraphs. If we have a set of safe path, or safe sequences, respectively, such that no two of them can appear on the same path of the DAG, then each of them must clearly appear in different paths of any path cover. Suppose that an ILP model has $k$ binary variables $x_{uvi}$, $i\in\{1,\dots,k\}$, for every arc $(u,v)$ of $G$, with the interpretation that $(u,v)$ appears in solution path $i$ iff $x_{uvi}=1$. Then, using such set of safe paths or sequences we can fix to 1 some of these binary variables, as in [acceleratingILP]. For example, we can assign the blue, orange, green and violet safe paths to solution paths 1,2,3 and 4, respectively (we show this assignment only for the blue and violet safe paths). We can proceed in a similar manner for safe sequences, now fixing more variables because sequences are overall longer (we show this assignment only for the blue and violet safe sequences).
@@
While safe paths fully capture contiguous safety information, it may be that the only way to reach a safe path $P_{2}$ is by passing through a safe path $P_{1}$, with some “complex” subgraph between $P_{1}$ and $P_{2}$. As such, the sequence $P_{1},P_{2}$ appears in some path of any path cover. Therefore, inspired by the results of Ma, Zheng and Kingsford [and_or_quant], we generalize the previous notion of safe paths to safe sequences of arcs that appears in all paths covers of a DAG $G$ (see Figure 1(b) for an example). More specifically, an application of the AND-Quant problem from [and_or_quant] is to characterize when a sequence $(e_{1},\dots,e_{t})$ of arcs of $G$ appears in all flow decompositions of $G$. Namely, when it holds that for any flow decomposition, there is a path $P_{i}$ of the decomposition such that $e_{1},\dots,e_{t}$ appear in $P_{i}$, in this order. This is based on solving a max flow problem from $(e_{1},\dots,e_{t})$ and $G$. In this paper we characterize safe sequences with respect to path covers of $G$. For this, we generalize our results for safe paths by using the notion of $u$-$v$ bridges (i.e. arcs belonging to all $u$-$v$ paths), for a suitable choice of nodes $u$ and $v$.
@@
[Maximal safe sequences enumeration]theoremmaxsafeseqsenum Given a DAG $G$ with $m$ arcs, there is an $O(m+o^{2})$-time algorithm computing all the maximal safe sequences of $G$, where $o$ is the total length of the output, namely of all maximal safe sequences.
@@ 2 Safe paths for path covers in $s$-$t$ DAGs
To avoid computing duplicated and non maximal safe paths we use the notion of core from [schmidt2023omnitig]. A core is a unitig such that its univocal extension results in a maximal safe path. Then, we contract all unitigs into an arc to obtain a contracted graph; clearly, there is a bijection between safe paths on the contracted graph and safe paths on the original graph. To compute the set of maximal safe paths we perform univocal extensions in every arc of the contracted graph whose tail has outdegree at least two (or is the source), and whose head has indegree at least two (or is the sink), while keeping a left and right pointer in the original graph that mimic the extensions on the contracted graph: initially, the left pointer points to the first node of the core and the right pointer points to its last node, and for every arc $a$ that we extend to in the contracted graph (to the left or right) the corresponding pointer extends over the whole unitig in the original graph corresponding to $a$. We refer to [schmidt2023omnitig] for the proofs of correctness of this procedure. The contraction step can be done in $O(m+n)$ time and the univocal extension phase can be done in $O(m+o)$ time where $o$ is the total length of the maximal safe paths, all requiring only linear space in the size of the graph. This algorithm solves Problem 1 in optimal time and space.
@@ 3 Safe sequences for path covers in $s$-$t$ DAGs
Let $h=(u,v)$ be an arc, $A$ the sequence of $s$-$u$ bridges, and $B$ the sequence of $v$-$t$ bridges. We argue that the concatenation of these sequences ordered by reachability is a safe sequence. First, it is a sequence, as the last arc of $A$ (if any) reaches $h$ and $h$ reaches the first arc of $B$ (if any). Secondly, some $s$-$t$ path $q$ of any path cover must cover $h$, and since, by construction, any $v$-$t$ path contains $B$ and any $s$-$u$ path contains $A$, $q$ necessarily contains the sequences $A$, $h$, and $B$ ordered by reachability. This shows that bridge extensions are safe.
Furthermore, as a consequence of Section 3, if $A$ is a safe sequence then there is special path $p$ containing an arc $h=(u,v)$ separating the forks of $A$ to the left of the merges of $A$ in $p$. Now note that the sequence of $s$-$u$ bridges must contain the arcs of $A$ preceding $h$ in $p$, since otherwise there would be a merge in the nodes of $p$ preceding $h$, a contradiction. Analogously, the sequence of $v$-$t$ bridges contains the arcs of $A$ succeeding $h$ in $p$, since otherwise there would be a fork in the nodes of $p$ succeeding $h$, a contradiction. Therefore any safe sequence is either the bridge extension of some arc in the sequence or a subsequence of the extension. From this discussion we can also bound the number of maximal safe sequences.
@@
Bridge extensions have analogous limitations to univocal extensions. Indeed, applying bridge extensions to two distinct arcs may result in the same sequence, and moreover the resulting sequence may not even be maximal. Nonetheless, from the discussion above, by extending every arc through its bridges, left and right, we are bound to find all the maximal safe sequences in the graph. We now present a lemma exhibiting a useful property of bridge extensions that will simplify the rest of this section. See Appendix 0.A for the proof.
@@
lemmaextsubstring If $A$ is a non maximal safe sequence computed via bridge extensions then it is a prefix or suffix of a maximal safe sequence.
@@
Given a DAG $G$, there is an $O(m^{2})$ time algorithm that computes all the maximal safe sequences of $G$.
@@
For every arc $(u,v)\in E(G)$ compute the set of $s$-$u$ and $v$-$t$ bridges. This computation can be done in $O(m)$ time using two DFS executions using, e.g., the algorithm proposed in [CAIRO2021103]. To remove duplicated, prefix, and suffix sequences we use an Aho Corasick Trie [aho1975efficient] over the set of computed sequences. This filtering step takes linear time proportional to the total length of the sequences, which is clearly $O(m^{2})$.
A faster algorithm.
Italiano et al. in [italiano] proved the following result. Given a strongly connected directed graph $G$, in $O(m+n)$ time and $O(n)$ auxiliary space it is possible to efficiently answer the following query: given two nodes $u$ and $v$, report all arcs $a\in E(G)$ such that $u$ does not reach $v$ or $v$ does not reach $u$ in $G-a$. Following the terminology of [italiano], we will call such an arc $a$ a separating arc for $u,v$. Importantly, therein they show this can be done in linear time on the number of separating arcs. Then, we can use the techniques of [italiano] together with our approach in Theorem 3.1 to solve Problem 2.
@@
The time spent to compute the separating $s$-$u$ arcs and the separating $v$-$t$ arcs is exactly the length of the safe sequence being computed. Then we spend $O(o)$ time to compute all the maximal safe sequences. If a safe sequence is not maximal, then it is a prefix or suffix of some maximal safe sequence since it was computed via bridge extensions. In the worst case we compute all the prefixes and suffixes of that sequence, which costs quadratic time in its length, thus, by summing up over the lengths of all prefixes and suffixes of all the the maximal safe sequences we get $O(o^{2})$. To conclude, we remove duplicated, prefix, and suffix sequences with an Aho Corasick Trie [aho1975efficient], which requires linear time proportional to $O(o^{2})$.
The drawback of Section 1.0.2 with respect to its analogue Figure 1 for safe paths is that we perform extensions on every arc instead on only those identifying maximal safe sequences, thus spending time computing non maximal sequences. Consequently, we also require a postprocessing phase to remove them. As such, we leave as open problem whether one can compute them in optimal $O(m+o)$ time, where $o$ denotes the length of all maximal safe sequences. We remark that the main ingredient to achieve the optimal running time for the analogous problem for paths (Problem 1) is that we are able to pinpoint in constant time the cores of the graph by spending only a linear time in a preprocessing step, that of building the contracted graph. A corresponding technique in the context of safe sequences would immediately lead to an optimal time algorithm.
@@ Appendix 0.A Missing proofs
Since $A$ is computed via bridge extensions there is an arc $h_{A}\in A$ whose bridge extension is $A$, and since $B$ is maximally safe there is an arc $h_{B}\in B$ whose bridge extension is $B$. If $h_{A}=h_{B}$ then they have the same bridge extension, so $h_{A}\neq h_{B}$.
Using Section 3, let $p$ be a path where every fork of $B$ precedes every merge of $B$ in $p$. Since $A$ is a subsequence of $B$, a merge (fork) node in $R_{b_{i}b_{i+1}}$ (for any $i$) with respect to $A$ is also a merge (fork) node with respect to $B$, and thus every fork of $A$ in $p$ precedes every merge of $A$ in $p$ (note that $B$ has at least two arcs, so $R_{b_{i}b_{i+1}}$ is well defined). If $h_{B}$ precedes $h_{A}$ in $p$ then, by construction of bridge extensions, the right bridge extension of $h_{A}$ coincides with the right bridge extension of $B$ restricted to the arcs succeeding $h_{A}$, i.e., there are no gaps in this suffix of $A$ with respect to $B$. Moreover, if between $h_{B}$ and $h_{A}$ there is no merge node with respect to $A$ then the left bridge extension of $h_{A}$ spans over the arcs between $h_{B}$ and $h_{A}$, and thus the extension would continue through $h_{B}$, implying that they would have the same left bridge extension. Then there is a merge node between $h_{B}$ and $h_{A}$, and so let $v\in p$ be the rightmost such node. Again, by construction of bridge extensions, the left bridge extension of $h_{A}$ contains every arc of $B$ between $v$ and $\textsc{t}(h_{A})$ and stops at $v$. This subsequence concatenated with $h_{A}$ and this concatenated with the right bridge extension of $A$ results in $A$, and this in turn is a suffix of $B$. If $h_{A}$ precedes $h_{B}$ we can argue analogously to conclude that $A$ is a prefix of $B$.

### v2

@@ 1 Introduction
Figure 1: Example of safe paths and safe sequences for the $s$-$t$ path covers in a DAG $G$ with unique source $s$ and unique sink $t$. That is, for every path cover of $G$ (i.e. set of paths from $s$ to $t$ in $G$ such that every arc of $G$ appears in a path of the cover), there is a path in the cover containing the safe path or the safe sequence. We can see that safe sequences extend safe paths over complex subgraphs. If we have a set of safe path, or safe sequences, respectively, such that no two of them can appear on the same path of the DAG, then each of them must clearly appear in different paths of any path cover. Suppose that an ILP model has $k$ binary variables $x_{uvi}$, $i\in\{1,\dots,k\}$, for every arc $(u,v)$ of $G$, with the interpretation that $(u,v)$ appears in solution path $i$ iff $x_{uvi}=1$. Then, using such set of safe paths or sequences we can fix to 1 some of these binary variables, as in [acceleratingILP]. For example, we can assign the blue, orange, green and violet safe paths to solution paths 1,2,3 and 4, respectively (we show this assignment only for the blue and violet safe paths). We can proceed in a similar manner for safe sequences, now fixing more variables because sequences are overall longer (we show this assignment only for the blue and violet safe sequences)
@@
While safe paths fully capture contiguous safety information, it may be that the only way to reach a safe path $P_{2}$ is by passing through a safe path $P_{1}$, with some “complex” subgraph between $P_{1}$ and $P_{2}$. As such, the sequence $P_{1},P_{2}$ appears in some path of any path cover. Therefore, inspired by the results of Ma, Zheng and Kingsford [and_or_quant], we generalize the previous notion of safe paths to safe sequences of arcs that appears in all paths covers of a DAG $G$ (see Figure 1(b) for an example). More specifically, an application of the AND-Quant problem from [and_or_quant] is to characterize when a sequence $(e_{1},\dots,e_{t})$ of arcs of $G$ appears in all flow decompositions of $G$. Namely, when it holds that for any flow decomposition, there is a path $P_{i}$ of the decomposition such that $e_{1},\dots,e_{t}$ appear in $P_{i}$, in this order. This is based on solving a max flow problem from $(e_{1},\dots,e_{t})$ and $G$. In this paper we characterize safe sequences with respect to path covers of $G$. For this, we generalize our results for safe paths by using the notion of $u$-$v$ bridges.
@@
[Maximal safe sequences enumeration]theoremmaxsafeseqsenum Given a DAG $G$ with $m$ arcs, there is an $O(m+o^{3})$-time algorithm computing a set of safe sequences containing all the maximal safe sequences of $G$, where $o$ is the total length of the output, namely of all maximal safe sequences.
@@ 2 Safe paths for path covers in $s$-$t$ DAGs
To avoid computing duplicated and non maximal safe paths we use the notion of core from [schmidt2023omnitig]. A core is a unitig such that its univocal extension results in a maximal safe path. With this, we proceed as follows. We contract all unitigs into an arc to obtain a contracted graph; clearly, there is a bijection between safe paths on the contracted graph and safe paths on the original graph. To compute the set of maximal safe paths we perform univocal extensions in every arc of the contracted graph whose tail has outdegree at least two (or is the source) and whose head has indegree at least two (or is the sink), while keeping a left and right pointer in the original graph that mimic the extensions on the contracted graph: initially, the left pointer points to the first node of the core and the right pointer points to its last node, and for every arc $a$ that we extend to in the contracted graph (to the left or right) the corresponding pointer extends over the whole unitig in the original graph corresponding to $a$. We refer to [schmidt2023omnitig] for the proofs of correctness of this procedure. The contraction step can be done in $O(m+n)$ time and the univocal extension phase can be done in $O(m+o)$ time where $o$ is the total length of the maximal safe paths, all requiring only linear space in the size of the graph. This algorithm solves Problem 1 in optimal time and space.
@@ 3 Safe sequences for path covers in $s$-$t$ DAGs
Let $h=(u,v)$ be an arc, $A$ the sequence of $s$-$u$ bridges, and $B$ the sequence of $v$-$t$ bridges. We argue that the concatenation of these sequences ordered by reachability is a safe sequence. First, it is a sequence, as the last arc of $A$ (if any) reaches $h$ and $h$ reaches the first arc of $B$ (if any). Secondly, some $s$-$t$ path $q$ of any path cover must cover $h$, and since, by construction, any $v$-$t$ path contains $B$ and any $s$-$u$ path contains $A$, $q$ necessarily contains the sequence $AhB$. This shows that bridge extensions are safe.
Furthermore, as a consequence of Section 3, if $A$ is a safe sequence then there is special path $p$ containing an arc $h=(u,v)$ separating the forks from the merges of $A$ in $p$. Now, the sequence of $s$-$u$ bridges must contain the arcs of $A$ preceding $h$ in $p$, since otherwise there would be a merge in the nodes of $p$ preceding $h$, a contradiction. Analogously, the sequence of $v$-$t$ bridges contains the arcs of $A$ succeeding $h$ in $p$, since otherwise there would be a fork in the nodes of $p$ succeeding $h$, a contradiction. Therefore any safe sequence is either the bridge extension of some arc in the sequence (or a subsequence of the extension, as dropping arbitrary arcs from a safe sequence yields another safe sequence). From this discussion we can also bound the number of maximal safe sequences.
@@
Bridge extensions have analogous limitations to univocal extensions. Indeed, applying bridge extensions to two distinct arcs may result in the same sequence, and moreover the resulting sequence may not be maximal. Nonetheless, from the discussion above, by extending every arc through its bridges, left and right, we are bound to find all the maximal safe sequences in the graph. We now present a lemma that will simplify the rest of this section. See Appendix 0.A for the proof.
@@
lemmaextsubstring If $A$ is a non maximal safe sequence computed via bridge extensions then it is a prefix, suffix, or prefix and suffix of a maximal safe sequence.
@@
Given a DAG $G$ with $m$ arcs, there is an $O(m^{2})$-time algorithm computing a set of safe sequences containing all the maximal safe sequences of $G$.
@@
For every arc $(u,v)\in E(G)$ compute the set of $s$-$u$ and $v$-$t$ bridges. This computation can be done in $O(m)$ time using two DFS executions using, e.g., the algorithm proposed in [CAIRO2021103].
A faster algorithm via dominators.
Another way of framing bridge extensions is via the well known notion of dominators. Given a digraph with a designated start vertex $s$, we say that an arc $(x,y)$ dominates an arc $(z,w)$ if all paths from $s$ to $z$ contain $a$. Clearly, $a$ is an $s$-$z$ bridge if and only if $a$ dominates $b$. There is an extensive study on dominators in the literature, namely that of computing the dominator tree of a general digraph. We refer to [chandak2016survey] for details.
Given a DAG $G=(V,E)$ and an arc $(u,v)\in E$, our goal compute is to compute $s$-$u$ bridges and $v$-$t$ bridges fast. We proceed as follows. First we build the dominator tree $T_{s}$ of $G$ rooted at $s$ and the dominator tree $T_{t}$ of $G$ with all its arcs reversed rooted at $t$. Now, left bridge extensions from an arc $a$ coincide with the traversal from $a\in V(T_{s})$ to $s$ in $T_{s}$. Analogously, the right bridge extensions from $a$ coincide with the traversal from $a\in V(T_{t})$ to $t$ in $T_{t}$. The correctness of the former is clear, and of the former comes from the fact that an arc $b$ is a $v$-$t$ bridge in the original graph if and only if $b$ is a $t$-$v$ bridge in the reversed graph.
A dominators’ related technique that we can use to compute bridge extensions is that of Italiano et al. in [italiano], where they showed the following result. Given a strongly connected directed graph $G$, in $O(m+n)$ time and $O(n)$ auxiliary space it is possible to efficiently answer the following query: given two nodes $u$ and $v$, report all arcs $a\in E(G)$ such that $u$ does not reach $v$ or $v$ does not reach $u$ in $G-a$. Following the terminology of [italiano], we will call such an arc $a$ a separating arc for $u,v$. Importantly, therein they show this can be done in linear time on the number of separating arcs. Then, we can use the techniques of [italiano] together to solve Problem 2.
@@
We analyze separately the time required to compute maximal and non maximal safe sequences. Note that the time spent to compute the separating $s$-$u$ arcs and the separating $v$-$t$ arcs is exactly the length of the safe sequence being computed. Let $s_{1}+s_{2}+\dots+s_{\ell}=o$, where $s_{i}$ denote the length of the $i$-th maximal safe sequence and $\ell\leq m$ by Corollary 3.
For maximal sequences we spend at least $O(o)$ time. In the worst case, each arc of every maximal safe sequence is a core for that sequence, and so we spend at most $s_{1}^{2}+s_{2}^{2}+\dots s_{\ell}^{2}=O(o^{2})$ time to compute all the duplicates of the maximal safe sequences.
By Corollary 3, each extension of an arc that resulted in a non maximal safe sequence is either a prefix, suffix, or prefix and suffix of a maximal safe sequence. In the worst case, every non maximal sequence is a prefix and suffix, as the sum of lengths of all prefixes (or suffixes) of a string of length $k$ is $O(k^{2})$ and that of all strings that are prefixes and suffixes is $O(k^{3})$. Then summing over all safe sequences that are prefixes and suffixes of maximal safe sequences we get $s_{1}^{3}+s_{2}^{3}+\dots+s_{\ell}^{3}=O(o^{3})$.
The drawback of Section 1.0.2 with respect to its analogue Figure 1 for safe paths is that we perform extensions on every arc, instead on only those identifying maximal safe sequences, i.e., the cores. As such, we leave as open problem whether one can compute the maximal sequences in optimal $O(m+o)$ time, where $o$ denotes the length of all maximal safe sequences. We remark that the main ingredient to achieve the optimal running time for the analogous problem for paths (Problem 1) is that we can pinpoint in constant time the cores of the graph by spending only a linear time in a preprocessing step, that of building the contracted graph. A corresponding technique in the context of safe sequences would immediately lead to an optimal time algorithm. Finally, we remark that any technique to build dominator trees fast can be used in our problem, namely the classical results presented in [tarjan1974finding, lengauer1979fast].
@@ Appendix 0.A Missing proofs
Let $B$ be a maximal safe sequence containing $A$. Since $A$ is computed via bridge extensions there is an arc $h_{A}\in A$ whose bridge extension is $A$, and since $B$ is maximally safe there is an arc $h_{B}\in B$ whose bridge extension is $B$. If $h_{A}=h_{B}$ then they have the same bridge extension, so $h_{A}\neq h_{B}$.
Using Section 3, let $p$ be a path where every fork of $B$ precedes every merge of $B$ in $p$. Since $A$ is a subsequence of $B$, a merge (fork) node in $R_{b_{i}b_{i+1}}$ (for any $i$) with respect to $A$ is also a merge (fork) node with respect to $B$, and thus every fork of $A$ in $p$ precedes every merge of $A$ in $p$ (note that $B$ has at least two arcs, so $R_{b_{i}b_{i+1}}$ is well defined). If $h_{B}$ precedes $h_{A}$ in $p$ then, by construction of bridge extensions, the right bridge extension of $h_{A}$ coincides with the right bridge extension of $B$ restricted to the arcs succeeding $h_{A}$, i.e., there are no gaps in this suffix of $A$ with respect to $B$. Moreover, if between $h_{B}$ and $h_{A}$ there is no merge node with respect to $A$ then the left bridge extension of $h_{A}$ spans over the arcs between $h_{B}$ and $h_{A}$, and thus the extension would continue through $h_{B}$, implying that they would have the same left bridge extension. Then there is a merge node between $h_{B}$ and $h_{A}$, and so let $v\in p$ be the rightmost such node. Again, by construction of bridge extensions, the left bridge extension of $h_{A}$ contains every arc of $B$ between $v$ and $\textsc{t}(h_{A})$ and stops at $v$. If there is an $s$-$v$ with no internal node in common with $p\setminus\{s\}$, we have that the subsequence between $v$ and $\textsc{t}(h_{A})$ concatenated with $h_{A}$ and this concatenated with the right bridge extension of $A$ results in $A$, and this in turn is a suffix of $B$. If the path witnessing the fact that $v$ is a merge intersects $p\setminus\{s\}$, then by the definition of bridge extension every arc of $B$ preceding the first node of this path will also be in $A$, and thus $A$ is a prefix and suffix of $B$. If $h_{A}$ precedes $h_{B}$ we can argue analogously to conclude that $A$ is a prefix or a prefix and suffix of $B$.